OpenStudy (anonymous):

Age Ticket Price Under 3 years free 3 to 12 years $35 Over 12 years $46 On a certain day, the company took 88 people on whale watching trips. There were 44 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under 3 years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

answer choices: (44 + x)35 + (88 - (44 + x)) ⋅ 46 = C (88 - x)35 + (88 - 44) ⋅ 46 = C (44 - x)35 + (88 - 44) ⋅ 46 = C (88 + x)35 + (88 - (44 + x)) ⋅ 46 = C

OpenStudy (anonymous):

the third equation

OpenStudy (anonymous):

@Sweetangel14 Can you maybe explain to me how you got that?

OpenStudy (anonymous):

x is the number of children under 3, so you have to subtract those from the 44 children 12 and under since they get in for free so you get 44-x , then you have to times whatever number to the ticket price for 3 to 12 yr olds so your equation is now (44-x)35, plus now the over 12s since you know there were 88 ppl and 44 of them were under 12, so you subtract 88-44 and then you times it by their ticket price so now the full equation is (44-x)35+(88-44)46=c